What is the derivative of cos4x


1
Expert's answer
2021-05-26T02:30:49-0400

Solution:

Let "y=\cos 4x"

Differentiating both sides w.r.t "x",

"y'=-\sin 4x \times (4x)'
=-4\sin 4x"